# Manifold Learning via Foliations and Knowledge Transfer

### v1

@@ Abstract
Understanding how real data is distributed in high dimensional spaces is the key to many tasks in machine learning. We want to provide a natural geometric structure on the space of data employing a deep ReLU neural network trained as a classifier. Through the data information matrix (DIM), a variation of the Fisher information matrix, the model will discern a singular foliation structure on the space of data. We show that the singular points of such foliation are contained in a measure zero set, and that a local regular foliation exists almost everywhere. Experiments show that the data is correlated with leaves of such foliation. Moreover we show the potential of our approach for knowledge transfer by analyzing the spectrum of the DIM to measure distances between datasets.
@@ 1 Introduction
The concept of manifold learning lies at the very heart of the dimensionality reduction question, and it is based on the assumption that we have a natural Riemannian manifold structure on the space of data [40, 19, 41, 14]. Indeed, with such assumption, many geometrical tools become readily available for machine learning questions, especially in connection with the problem of knowledge transfer [9, 45], as geodesics, connections, Ricci curvature and similar [1]. In particular, the fast developing field of Information Geometry [4, 27], is now providing with the techniques to correctly address such questions [26].
However, the practical situation, arising for example in classifying benchmark datasets as MNIST [23], Fashion-MNIST [49] and similar, with CNNs (Convolutional Neural Networks), is generally more complicated and does not allow for such simple description, at least in the cases most relevant for concrete applications, and calls for more sophisticated mathematical modeling.
@@
Machine learning models can be roughly divided into two categories: classifier and generative models. In particular, Deep Learning vision classifier models, via their intrinsic hierarchical structure, offer a natural representation and implicit organization of the input data [31]. For example, the authors of [18] show how a multilayer encoder network can be successfully employed to transform high dimensional data into low dimensional code, then to retrieve it via a “decoder”.
Such organisation strikingly mirrors the one occurring in the human visual perception, as also observed in the seminal work [41], and later an inspiration for the spectacular successful employment of CNNs for supervised classification tasks [24].
@@
In this paper, we want to provide the data space of a given dataset with a natural geometrical structure, and then employ such structure to extract key information. We will employ a suitably trained CNN model to discern a foliation structure in the data space and show experimentally how the dataset our model has been trained with is strongly correlated with its leaves, which are submanifolds of the data space. The mathematical idea of foliation is quite old (see [33, 13] and refs therein); however, its applications in control theory (see [3] and refs therein) via sub-Riemannian geometry and machine learning have only recently become increasingly important [42, 15].
@@
The foliation structure on the data space, discerned by our model, however, is non-standard and presents singular points. These are points admitting a neighbourhood where the rank of the distribution, tangent at each point to a leaf of the foliation, changes. Moreover, in the presence of typical non linearities of the network, as ReLU, there are also non smooth points, thus hindering in such points the smooth manifold structure itself of the leaf. However, we prove that both singular and smooth points are a measure zero set in the data space, so that the foliation is almost everywhere regular and its distribution well defined. As it turns out in our experiments, the samples belonging to the dataset we train our model with are averagely close to the set of singular points. It forces us to model the data space with a singular foliation that we call learning foliation in analogy with the learning manifold. Applications of singular foliations were introduced in connection with control theory [38]; their study is currently an active area of investigation [22]. As we shall show in our experiments, together with their distributions, singular foliations provide with an effective tool to single out the samples belonging to the datasets the model was trained with and at the same time discern a notion of distance between different datasets belonging to the same data space.
@@
Our paper is organized as follows. In Sec. 2, we recap the previous literature, closely related to our work. In Sec. 3, we start by recalling some known facts about Information Geometry in 3.1. Then, in Sec. 3.2 we introduce the data information matrix (DIM), defined via a given model, and two distributions $\mathcal{D}$ and $\mathcal{D}^{\perp}$ naturally associated with it, together with some illustrative examples elucidating the associated foliations and their leaves. In Sec. 3.3 we introduce singular distributions and foliations and then we prove our main theoretical results expressed in Lemma 3.4 and Theorem 3.6. Lemma 3.4 studies the singularities of the distribution $\mathcal{D}$. Theorem 3.6 establishes that the singular points for $\mathcal{D}$ are a measure zero set in the data space. Hence our foliation, though singular, acquires geometric significance. Finally in Sec. 4 we elucidate our main results with experiments. Moreover, we show how the foliation structure and its singularities can be exploited to determine which dataset the model was trained on, and its distance from similar datasets in the same data space. We also make some "proof of concept" considerations regarding knowledge transfer to show the potential of the mathematical singular local foliation and distribution structures in data space.
@@ 2 Related Works
The use of Deep Neural Networks as tools for both manifold learning and knowledge transfer has been extensively investigated. The question of finding a low dimensional manifold structure (latent manifold) into high dimensional dataset spaces starts with PCA and similar methods to reach the more sophisticated techniques as in [14, 26], (see also [29] for a description of the most popular techniques in dimensionality reduction and [10] for a complete bibliography or the origins of the subject). Such understanding was applied towards the knowledge transfer questions in [11, 7] and more recently [12]. Though this is not our approach, it is important to mention that one important technique of knowledge transfer is via shared parameters as in [28]. The idea of using techniques of Information Geometry for machine learning started with [4]. More recently, it was used for manifold learning in [26] (see also refs therein). Employing foliations to reduce dimensionality is not per se novel. In [15], the authors introduce some foliation structure in the data space of a neural network, with the aim of approaching dimensionality reduction. In [42], orthogonal foliations in the data space are used to create adversarial attacks, and provide the data space with curvature and other Riemannian notions to analyse such attacks. In [39] invariant foliations are employed to produce a reduced order model for autoencoders. Also, in considering singular points for our foliation, we are led to study the singularities of a neural network. This was investigated, for the different purpose of network expressibility in [16].
3 Methodology
3.1 Information Geometry
The statistical simplex consists of all probability distributions in the form $p(y|x,w)=(p_{1}(y|x,w),\dots,p_{c}(y|x,w))$, where $x$ is a data point, belonging to a certain dataset $D\subset\mathbf{R}^{d}$, divided into $c$ classes, while $w\in\mathbf{R}^{N}$ are the learning parameters [4, 20, 27]. We define the information loss as: $I(x,w)=-\log(p(y|x,w))$ and the Fisher information matrix (FIM) [32] as
$F(x,w)=\mathbf{E}_{y\sim p}[\nabla_{w}\log p(y|x,w)\cdot(\nabla_{w}\log p(y|x,w))^{T}]$ (1)
@@ 3.1 Information Geometry
$D(x,w)=\mathbf{E}_{y\sim p}[\nabla_{x}\log p(y|x,w)\cdot(\nabla_{x}\log p(y|x,w))^{T}]$ (2)
As noted in [26], some directions in the parameter or data space, may be more relevant than others in manifold learning theory. We have the following result [15], obtained with a direct calculation.
@@
The Fisher information matrix $F(x,w)$ and the data information matrix $D(x,w)$ are positive semidefinite symmetric matrices. Moreover:
$\begin{array}[]{c}\ker F(x,w)=(\mathrm{span}_{i=1,\ldots,c}\{\nabla_{w}\log p_{i}(y|x,w)\})^{\perp},\\ \\ \ker D(x,w)=(\mathrm{span}_{i=1,\ldots,c}\{\nabla_{x}\log p_{i}(y|x,w)\})^{\perp}.\end{array}$ (3)
where the orthogonal is taken with respect to the euclidean product. In particular $\mathrm{rank}\ F(x,w)<c$ and $\mathrm{rank}\ D(x,w)<c$.
@@
The result (3) suggests to consider the distribution $\mathcal{D}$:
$\begin{array}[]{c}\mathbf{R}^{d}\ni x\mapsto\mathcal{D}_{x}:=\mathrm{span}_{i=1,\ldots,c}\{\nabla_{x}\log p_{i}(y|x,w)\}\subset T_{x}\mathbf{R}^{d}\end{array}$ (4)
where we assume the learning parameters $w\in\mathbf{R}^{N}$ to be fixed as $x$ varies. In general a distribution on a manifold $M$ assigns to each point a subspace of the tangent space to $M$ at that point (see [43, 44] for more details), in our case $M=\mathbf{R}^{d}$.
@@ 3.2 Distributions and Foliations
Whenever we have a distribution on a manifold $M$, it is natural to ask whether it is integrable. A distribution $\mathcal{C}$ on $M$ is integrable if for every $x\in M$, there exists a connected immersed submanifold $N$ of $M$, with $T_{y}N=\mathcal{C}_{y}$ for all $y\in N$. In other words, the distribution defines at each point the tangent space to a submanifold. Whenever a distribution is integrable, we have a foliation, that is, $M$ becomes a disjoint union of embedded submanifolds called the leaves of the foliation. If $\dim\mathcal{C}_{x}=\dim\mathcal{C}_{y}$ for all $x,y\in M$ we say that the foliation, or the corresponding distribution, has constant rank. Fig. 1 illustrates a constant rank foliation and its orthogonal complement foliation, with respect to the euclidean metric in $M=\mathbf{R}^{3}$ (the ambient space).
$\mathbf{R}^{d}$
$\mathcal{L}_{\mathcal{D}^{\perp}}$
$\mathcal{L}_{\mathcal{D}}$
Figure 1: Foliations: $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{D}^{\perp}}$ denote the set of the leaves in distributions $\mathcal{D}$ and $\mathcal{D}^{\perp}$ respectively.
Under suitable regularity hypothesis, Frobenius theorem [43, 44] provides with a characterization of integrable distributions.
Theorem 3.2.
(Frobenius Theorem). A smooth constant rank distribution $\mathcal{C}$ on a real manifold $M$ is integrable if and only if it is involutive, that is for all vector fields $X,Y\in\mathcal{C}$, $X,Y]\in\mathcal{C}$.  [^1] If $X$, $Y$ are vector fields on a manifold, we define their Lie bracket as [X,Y]:=XYYX[X,Y]:=XY-YX.
To elucidate this result in Fig. 4 we look at two examples of foliations of 1-dimensional distributions, obtained from the distribution $\mathcal{D}$ in (4) of a neural network trained on the Xor function, with non linearities GeLU and ReLU [^2] We look only at smooth points, which form an open set.. Notice that since in both cases $\mathcal{D}$ is 1-dimensional, then it is automatically integrable (bracket of vector fields being zero). We have the following result; the proof is a simple calculation based on the Frobenius Theorem [15].
@@
Let the notation be as above. For an empirical probability $p$ defined via a deep ReLU neural network the distribution $\mathcal{D}$ in the data space:
$x\mapsto\mathcal{D}_{x}=\mathrm{span}_{i=1,\ldots,c}\{\nabla_{x}\log p_{i}(y|x,w)\}$
@@
We call the foliation resulting from Prop. 3.3 the learning foliation. Its significance is clarified by the following examples.
Figure 2: Moving according to $\mathcal{D}$.
Figure 3: Moving according to $\mathcal{D}^{\perp}$.
As an illustration of the distributions $\mathcal{D}$ and $\mathcal{D}^{\perp}$ (orthogonal computed with respect to the euclidean metric in $M=\mathbf{R}^{d}$), in MNIST, we present Fig. 2 and Fig. 3. We notice that, while moving in $\mathcal{D}$ from one data point , the model predicts a meaningful label, while moving in $\mathcal{D}^{\perp}$, the model maintains high confidence on the label, though not in line with the image. Both Fig. 2 and Fig. 3 were obtained by projecting the same direction on $\mathcal{D}$ and $\mathcal{D}^{\perp}$ at each step. Notice that the phenomena we observe above are due to the fact that the neural network output probabilities are invariant by moving in the kernel of the DIM, that is the distribution $\mathcal{D}^{\perp}$.
For a model with GeLU non linearity however, one can see experimentally that we do not have the involutivity property anymore for the distribution $\mathcal{D}$ (see Table 2). Hence, there is no foliation whose leaves fill the data space, naturally associated to it via Frobenius Theorem. To see this more in detail, we define the space $\mathcal{V}^{D}_{x}$ generated by $\mathcal{D}_{x}$ and the Lie brackets of their generators:
@@
Table 1: Involutivity of the distribution $\mathcal{D}$
@@
In Table 2 we report averages of the dimensions of the spaces $\mathcal{D}_{x}$, $\mathcal{V}^{D}_{x}$ for a sample of 100 points $x\in\mathbf{R}^{d}$. The non involutivity of the distribution is deduced from the fact the dimension increases when we take the space spanned by the distribution and the brackets of its generators. As we can see, while for the ReLU non linearity $\mathcal{D}$ is involutive and we can define the learning foliation, the brackets of vector fields generating $\mathcal{D}$ do not lie in $\mathcal{D}$ for the GeLU and sigmoid non linearities. Consequently, there is no foliation and the sub Riemannian formalism appears more suitable to describe the geometry in this case. We shall not address the question here.
3.3 Singular Foliations

### v2

@@ Abstract
Understanding how real data is distributed in high dimensional spaces is the key to many tasks in machine learning. We want to provide a natural geometric structure on the space of data employing a ReLU neural network trained as a classifier. Through the Data Information Matrix (DIM), a variation of the Fisher information matrix, the model will discern a singular foliation structure on the space of data. We show that the singular points of such foliation are contained in a measure zero set, and that a local regular foliation exists almost everywhere. Experiments show that the data is correlated with leaves of such foliation. Moreover we show the potential of our approach for knowledge transfer by analyzing the spectrum of the DIM to measure distances between datasets.
keywords:
ReLU Networks, Deep Learning, Data Information Matrix, Singular Foliations, Learning Manifolds, Dataset Distances, Knowledge Transfer, Information Geometry.
@@ 1 Introduction
The concept of manifold learning lies at the very heart of the dimensionality reduction question, and it is based on the assumption that there exists a natural Riemannian manifold structure on the space of data [ten1998, hinton2003, ten2000, feff2016]. Indeed, with such assumption, many geometrical tools become readily available for machine learning questions, especially in connection with the problem of knowledge transfer [bozi2020, weiss2016], as geodesics, connections, Ricci curvature and similar [ache2018]. In particular, the fast developing field of Information Geometry [amari1998, martens2020], is now providing with the techniques to correctly address such questions [maillet].
However, the practical situation that arises, for instance, in the classification task of benchmark datasets such as MNIST [lecun1998], Fashion-MNIST [fashionmnist] and similar ones is usually too high-dimensional to allow for such straightforward description. This complexity requires for more sophisticated mathematical modelling, which we shall explore here by replacing the notion of manifold with that of singular foliation.
@@
Deep Learning vision classifier models, via their intrinsic hierarchical structure, offer a natural representation and implicit organization of the input data [olah]. For example, the authors in [hinton2006] show how a multilayer encoder network can be successfully employed to transform high dimensional data into low dimensional code, then to retrieve it via a “decoder”.
@@
In this paper, we want to provide the data space of a given dataset with a natural geometrical structure, and then employ such structure to extract key information. We will employ a suitably trained neural network model to define a foliation structure in the data space and show experimentally how the training set of our model is strongly correlated with its leaves, which are submanifolds of the data space. The mathematical idea of foliation is quite old (see [reeb, ehre] and refs therein); however, its applications in control theory (see [agra2004] and refs therein) via sub-Riemannian geometry and machine learning have only recently become increasingly important [tron2022, gf2021].
@@
The foliation structure on the data space, discerned by our model, however, is non-standard and presents singular points. These are points admitting a neighbourhood where the rank of the distribution, tangent at each point to a leaf of the foliation, changes. Moreover, in the presence of typical non-smooth activation functions of the network (e.g. ReLU, LeakyReLU, Maxpool or similar), there are also non-smooth points where the manifold structure of the leaf itself is questioned. However, we prove that both singular and non-smooth points are a measure zero set in the data space, so that the foliation is almost everywhere regular and its distribution well defined. As it turns out in our experiments, the samples belonging to the dataset we train our model with are averagely close to the set of singular points. It forces us to model the data space with a singular foliation that we call data foliation in analogy with the data manifold. Applications of singular foliations were introduced in connection with control theory [suss1973]; their study is currently an active area of investigation [lavau2017]. As we shall show in our experiments, singular foliations and distributions provide with an effective tool to single out the samples belonging to the datasets the model was trained with and at the same time discern a notion of distance between different datasets belonging to the same data space.
Main contributions.
  • The introduction of a singular geometrical framework to study neural networks with the Data Information Matrix (DIM).
  • The proof of its well-definednes (a.e.) for ReLU neural networks, given by theorem˜3.6.
  • The experimental study of the DIM’s eigenvalues and rank on different datasets, which sheds light on its understanding of the network’s training.
  • A toy example (knowledge transfer) that provides a perspective on the relevance of our framework for concrete applications in machine learning.
Organization of the paper.
In section˜2, we recap the previous literature, closely related to our work.
@@
In section˜3, we start by recalling some known facts about Information Geometry in section˜3.1. Then, in section˜3.2 we introduce the data information matrix (DIM), defined via a given model, and two distributions $\mathcal{D}$ and $\mathcal{D}^{\perp}$ naturally associated with it, together with some illustrative examples elucidating the associated foliations and their leaves. In section˜3.3 we introduce singular distributions and foliations and then we prove our main theoretical results expressed in lemma˜3.4 and theorem˜3.6. lemma˜3.4 studies the singularities of the distribution $\mathcal{D}$. theorem˜3.6 establishes that the singular points for $\mathcal{D}$ are a measure zero set in the data space. Hence our foliation, whose existence relies on mathematical results theorem˜3.2, that we apply as in proposition˜3.3, though singular, acquires geometric significance.
Finally in section˜4 we elucidate our main results with experiments. Moreover, we show how the foliation structure and its singularities can be exploited to determine which dataset the model was trained on, and its distance from similar datasets in the same data space. We also make some “proof of concept” considerations regarding knowledge transfer to show the potential of the mathematical singular local foliation and distribution structures in data space.
@@ 2 Related Works
The use of Neural Networks as tools for both manifold learning and knowledge transfer has been extensively investigated. We give a brief overview on the literature closer to our contribution.
Latent manifold. The question of finding a low dimensional manifold structure (latent manifold) in high dimensional dataset spaces starts with PCA and similar methods to reach the more sophisticated techniques as in [feff2016, maillet]. The reader can also look at [murphy2022] for a description of the most popular techniques in dimensionality reduction and [burges2010] for a complete bibliography or the origins of the subject. Such understanding was applied towards the knowledge transfer questions in [cook2007, ben2012] and more recently in [dikkala2021]. Moreover, distances between datasets were also studied through Optimal Transport in [Alvarez-Melis2019Geometric, Hua2023DynamicFO]. Another important technique of knowledge transfer is via shared parameters as in [maurer2016]. The idea of using techniques of Information Geometry for machine learning started with [amari1998]. More recently, it was used for manifold learning in [maillet] (see also refs therein).
Foliations. Employing foliations to reduce dimensionality is not per se novel. In [gf2021], the authors introduce some foliation structure in the data space of a neural network, with the aim of approaching dimensionality reduction exploiting the low rank of the equivalent of the Fisher matrix on the dataspace. However, they cannot conclude on the link between the manifold hypothesis and the considered foliation. Dimensionality reduction remains then a conjecture. Observations on the low rank of the Fisher matrix, and its key role in CNN parameter space, appear also in [Sun2017RelativeFI, Sun2024]. In [tron2022], orthogonal foliations in the data space are used to create adversarial attacks, and provide the data space with curvature and other Riemannian notions to analyze such attacks. In [szalai2023] invariant foliations are employed to produce a reduced order model for autoencoders. Also, in considering singular points for our foliation, we are led to study the singularities of a neural network. This was investigated, for the different purpose of network expressibility in [hanin2019].The partitioning of the data space into connected components after removing non-smooth points of ReLU CNNs has been studied in [Hanin2019ComplexityOL] and in [Montfar2014OnTN].
3 Foliations for ReLU Neural Networks
3.1 The Information Geometry of Neural Networks
In this article, we will focus our attention on neural networks $N_{w}$ with linear layers, ReLU activation functions and a Softmax at the end, that we call ReLU neural networks. In fact, the results of this article can easily be extended to all piecewise linear activation functions such as MaxPool or AveragePool. More precisely, the network is given by $N_{w}=\operatorname{Softmax}\circ S_{w}$ with the score function (or logits) $S_{w}$ being a composition of linear layers and activation functions defined by:
$S_{w}(x)=L_{W_{\ell}}\circ\sigma\circ\dots\circ\sigma\circ L_{W_{1}}$ (1)
where $\sigma$ is the ReLU non linearity, $L_{W_{i}}$ are linear layers (including bias), $w$ is the collection of all the parameters of the model (e.g. the $W_{i}$’s) and $\ell$ is the total number of linear layers.
The goal of this network will be the classification of the input $x\in\mathbf{R}^{d}$ between $c$ classes $\{y_{1},\ldots,y_{c}\}$. The function $N_{w}$ associates to each input $x$ the probability of belonging to each class, that we will denote by the vector
$p:=N_{w}(x)=\left(p\left(y_{1}\mid x,w\right),\ldots,p\left(y_{c}\mid x,w\right)\right)$ (2)
In these settings, the output of the network gives the parameters of a categorical distribution from which the predicted class $Y$ may be sampled from.
We define the information loss as: $I(x,w)=-\ln(p(y|x,w))$ and the Fisher information matrix (FIM) [rao1945] as
$F_{i,j}(x,w):=\mathbf{E}_{Y\mid x,w}\left[\left(\partial_{w^{i}}\ln p\left(Y\mid x,w\right)\right)\left(\partial_{w^{j}}\ln p\left(Y\mid x,w\right)\right)\right],\quad 1\leq i,j\leq N.$ (3)
@@ 3.1 The Information Geometry of Neural Networks
$D_{i,j}(x,w):=\mathbf{E}_{Y\mid x,w}\left[\left(\partial_{x^{i}}\ln p\left(Y\mid x,w\right)\right)\left(\partial_{x^{j}}\ln p\left(Y\mid x,w\right)\right)\right],\quad 1\leq i,j\leq d.$ (4)
To compute the DIM in practice, one must then compute the Jacobian matrix of the probability density $p$, i.e. the output of the network $N_{w}$, with respect to its input $x$. This is typically achieved through automatic differentiation [baydinAutomaticDifferentiationMachine2017], implemented in most machine learning libraries.
As noted in [maillet], some directions in the parameter or data space, may be more relevant than others in manifold learning theory. We have the following result [gf2021], obtained with a direct calculation.
@@
The Fisher information matrix $F(x,w)$ and the data information matrix $D(x,w)$ are positive semidefinite symmetric matrices. Moreover, if the $\mathrm{span}$ of a set of vectors denotes the set of all their linear combinations,
$\begin{array}[]{c}\ker F(x,w)=(\mathrm{span}_{i=1,\ldots,c}\{\nabla_{w}\ln p(y_{i}|x,w)\})^{\perp},\\ \\ \ker D(x,w)=(\mathrm{span}_{i=1,\ldots,c}\{\nabla_{x}\ln p(y_{i}|x,w)\})^{\perp},\end{array}$ (5)
where the orthogonal $\perp$ is taken with respect to the Euclidean product. In particular,
$\mathrm{rank}\ F(x,w)<c\quad\text{and}\quad\mathrm{rank}\ D(x,w)<c.$
@@
Recall that a distribution is the association to each point $x$ of a linear subspace of the tangent space at $x$. The result (5) then suggests to consider the distribution $\mathcal{D}$:
$\begin{array}[]{c}\mathbf{R}^{d}\ni x\mapsto\mathcal{D}_{x}:=\mathrm{span}_{i=1,\ldots,c}\{\nabla_{x}\ln p(y_{i}|x,w)\}\subset T_{x}\mathbf{R}^{d}\end{array}$ (6)
where we assume the learning parameters $w\in\mathbf{R}^{N}$ to be fixed as $x$ varies in the dataspace. In general a distribution on a manifold $M$ assigns to each point a subspace of the tangent space to $M$ at that point (see [tu2008, tu2017] for more details), in our case $M=\mathbf{R}^{d}$.
Observation 1.
[…]
@@ 3.2 Distributions and Foliations
Whenever we have a distribution on a manifold $M$, it is natural to ask whether it is integrable. A distribution $\mathcal{C}$ on $M$ is integrable if for every $x\in M$, there exists a connected immersed submanifold $N$ of $M$, with $T_{y}N=\mathcal{C}_{y}$ for all $y\in N$. In other words, the distribution defines at each point the tangent space to a submanifold. Whenever a distribution $\mathcal{C}$ is integrable, we have a foliation, denoted by $\mathcal{L}_{\mathcal{C}}$. The manifold $M$ becomes a disjoint union of embedded submanifolds called the leaves of the foliation. If $\dim\mathcal{C}_{x}=\dim\mathcal{C}_{y}$ for all $x,y\in M$ we say that the foliation, or the corresponding distribution, has constant rank. figure˜1 illustrates a constant rank foliation and its orthogonal complement foliation, with respect to the Euclidean metric in $M=\mathbf{R}^{3}$ (the ambient space).
Figure 1: Schematic illustration of foliations: $\mathcal{L}_{\mathcal{D}}$ and $\mathcal{L}_{\mathcal{D}^{\perp}}$ denote the set of the leaves in distributions $\mathcal{D}$ and $\mathcal{D}^{\perp}$ respectively. The images represent points in the data space when moving tangentially to $\mathcal{L}_{\mathcal{D}}$ (left) or to $\mathcal{L}_{\mathcal{D}^{\perp}}$ (bottom).
In this paper, we focus on the study of the distribution $\mathcal{D}$ defined in (6), strictly related with the DIM $D$, as defined in (4), since $\mathcal{D}$ is equivalently spanned by the columns of $D$ (see ˜1). Under suitable regularity hypothesis, Frobenius theorem [tu2008, tu2017] provides with a characterization of integrable distributions.
Theorem 3.2 (Frobenius Theorem).
A smooth constant rank distribution $\mathcal{C}$ on a real manifold $M$ is integrable if and only if it is involutive, that is for all vector fields $X,Y\in\mathcal{C}$, $X,Y]\in\mathcal{C}$.  [^1] If $X$, $Y$ are vector fields on a manifold, we define their Lie bracket as [X,Y](f):=X(Y(f))Y(X(f))[X,Y](f):=X(Y(f))-Y(X(f)) for every real function $f$.
[…]
(a) GeLU.
(b) ReLU.
Figure 2: The blue lines are a sample of the data foliation defined by the distribution $\mathcal{D}$ (6) for a Xor network. The two classes of the Xor problem are represented in red and green squares underneath. The red dot is a singular point for the foliation. In (b), the green lines are the non-smooth points.
In a more general setting, especially relevant for the applications, we have the following result in dimension $d$ with $c\geq 2$.
@@
Let the notation be as above. For an empirical probability $p$ defined by the output (normalized by softmax) of a ReLU neural network the distribution $\mathcal{D}$ in the data space:
$x\mapsto\mathcal{D}_{x}=\mathrm{span}_{i=1,\ldots,c}\{\nabla_{x}\ln p(y_{i}|x,w)\}$
@@
The proof of this proposition can be found in Appendix B. We call the foliation resulting from proposition˜3.3 the data foliation. Its significance is clarified by the following examples.
As an illustration of the distributions $\mathcal{D}$ and $\mathcal{D}^{\perp}$ (orthogonal computed with respect to the Euclidean metric in $M=\mathbf{R}^{d}$), in MNIST, we represent in figure˜1 pictures of points along paths tangential to $\mathcal{D}$ (left) or $\mathcal{D}^{\perp}$ (bottom). These pictures were obtained by projecting the same direction on $\mathcal{D}$ (left) or $\mathcal{D}^{\perp}$ (bottom) at each step. Note that moving along a path in $\mathcal{L}_{\mathcal{D}^{\perp}}$ leaves the predicted probability and label invariants, while moving along a path in $\mathcal{L}_{\mathcal{D}}$ changes the predicted label to adapt to the new picture.
For a model with GeLU non linearity however, one can see experimentally that we do not have the involutivity property anymore for the distribution $\mathcal{D}$ (see table˜1). This can be understood as coming from the non-vanishing Hessian matrix of the GeLU function, as detailed in Appendix B. Hence, there is no foliation whose leaves fill the data space, naturally associated to $\mathcal{D}$ via Frobenius Theorem. To see this more in detail, we define the space $\mathcal{V}^{D}_{x}$ generated by $\mathcal{D}_{x}$ and the Lie brackets of their generators:
@@
Table 1: Involutivity of the distribution $\mathcal{D}$: dimension averaged on 100 points sampled in $\left[0,1\right]^{d}$.
@@
In table˜1 we report averages of the dimensions of the spaces $\mathcal{D}_{x}$, $\mathcal{V}^{D}_{x}$ for a sample of 100 points $x\in\left[0,1\right]^{d}$ with a network trained on MNIST. The non involutivity of the distribution is deduced from the fact that the dimension increases when we take the space spanned by the distribution and the brackets of its generators. As we can see, while for the ReLU non linearity $\mathcal{D}$ is involutive and we can define the data foliation, the brackets of vector fields generating $\mathcal{D}$ do not lie in $\mathcal{D}$ for the GeLU and sigmoid non linearities. Consequently, there is no foliation and the sub Riemannian formalism appears more suitable to describe the geometry in this case. We shall not address the question here.
3.3 Singular Foliations and Main Result
In this section, we give our main result namely theorem˜3.6, regarding the singular points of $\mathcal{D}$.